# LoC-Path: Learning to Compress for Pathology Multimodal Large Language Models

### v1

@@ 3 Methodology
A WSI is decomposed into a long sequence of tiles. For each training image-text pair, we denote the tile-level tokens from the tile-level encoder by $\bm{X}=[\bm{x}_{1},\ldots,\bm{x}_{N}]\in\mathbb{R}^{N\times D_{v}}$, with 2D-grid coordinates $\bm{C}=(c_{i},r_{i})\in\mathbb{N}^{2}$, where $N$ is the number of the tile-level tokens, $D_{v}$ is the tile-level encoder’s feature dimension. $c_{i}$ and $r_{i}$ are the column and row grid coordinates with stride $1$, respectively. In MAE pre-training task, we produce $L$-length visual latents $\bm{Z}\in\mathbb{R}^{L\times D_{v}^{\prime}}$, where $D_{v}^{\prime}$ is the embedding dimension of MAE resampler. The text query tokens are $\bm{Q}_{\text{text}}$ and the answer labels are $\bm{Y}$. In single-round visual question answering (VQA) tasks, $\bm{\hat{Y}}$ is generated by an MLLM for a given $\bm{X}$ and text query $\bm{Q}_{\text{text}}$.
@@ 6 Redundancy and Task Relevance Analysis
$F(p)=\frac{\sum_{i=1}^{\lfloor pN\rfloor}\max(s_{i},0)}{\sum_{i=1}^{N}\max(s_{i},0)},\qquad p\in[0,1].$
@@ 7 Pairwise Ranking Loss Details
To compute this loss, we first sort the Top-$M$ indices $I$ according to the teacher distribution $\bm{q}$, yielding $I_{\text{sorted}}$. We select $n=\min(M/2,N_{\text{pairs}})$ pairs, where $N_{\text{pairs}}$ is a hyperparameter (e.g., 8). We define the ”high” set $I_{\mathrm{hi}}=I_{\text{sorted}}[:n]$ (the top $n$ indices) and the ”low” set $I_{\mathrm{lo}}=I_{\text{sorted}}[M//2:M//2+n]$ (the first $n$ indices from the bottom half).
@@ 8 Implementation and Training Details
Training Details. The resampler is pre-trained for $100\mathrm{k}$ iterations on CONCH-V1 [32] WSI features with a mini-batch size of $2$ per GPU and gradient accumulation $4$ on $4$ RTX A6000 Ada GPUs (effective global batch size $32$). We randomly mask $75\%$ of the input tokens and optimize the MAE objective $\mathcal{L}_{\mathrm{MAE}}$ introduced in Sec. 3.1, with $\lambda_{\mathrm{cover}}=5\times 10^{-4}$ for the coverage term $\mathcal{L}_{\mathrm{cover}}$ and $\lambda_{\mathrm{feat}}=1\times 10^{-3}$ for the diversity term $\mathcal{L}_{\mathrm{feat}}$. We use AdamW with an initial learning rate $1\times 10^{-4}$, a $3\%$ warm-up phase, and a cosine learning-rate schedule. During this pre-training we gradually anneal the Top-$K$ in the resampler cross-attention from $(128,64)$ to $(64,32)$ following a cosine schedule.
We then train the first fusion stage for $3$ epochs with a global batch size of $512$ (batch size $2$ per GPU and gradient accumulation $64$ on $4$ GPUs) and an initial learning rate $1\times 10^{-4}$. In this stage we freeze both the Qwen2.5-7B-Instruct LLM and the resampler, and only train the visual projector together with the CARA+TIS modules on top of the LLM. The training objective is the multimodal loss as defined in Sec. 3.2, where we set $\lambda_{\mathrm{TIS}}=\lambda_{\mathrm{rank}}=0.02$ for the self-distillation (TIS) loss $\mathcal{L}_{\mathrm{TIS}}$ and the ranking loss $\mathcal{L}_{\mathrm{rank}}$.
In the second fusion stage, we unfreeze the resampler’s last cross-attention layer and its latent in-projection, and insert LoRA modules (rank $128$, $\alpha=64$) into all linear layers of the LLM. We then train all unfrozen components on all VQA pairs in WSI-Bench for $2$ epochs with a global batch size of $128$ and an initial learning rate $2\times 10^{-5}$, again using the same multimodal fusion loss with $\lambda_{\mathrm{TIS}}=\lambda_{\mathrm{rank}}=0.02$. Across all training stages, we use Xtuner [10] with DeepSpeed ZeRO-2 [40] and FlashAttention 2 [11] for efficient training; FA2 is applied to the LLM self-attention but not to the cross-attention layers. A cosine learning-rate scheduler is used in all stages, and all experiments are conducted on $4$ RTX A6000 Ada GPUs.
Training Optimization. The pre-training requires about $10$ hours. The first fusion stage takes about $1$ hour. The second fusion stage requires about $8$ hours to $9$ hours.
@@
With file format conversion, we are able to load long sequences with very low I/O costs. Unlike the conventional tile-level data loader in digital pathology [9, 51], loading one data point in a single batch, we are able to load multiple tile-level sequences in a single batch simultaneously. To deal with sequences with different lengths, we pack these sequences into the same length (we pad sequences with different lengths to the same $60k$ lengths and load padded sequences with their corresponding padding masks). As a result, we can increase the batch size to more than one (e.g 4) on each GPU. Doing so allows us to reduce the data loading rounds if we set the max training epochs to be $2$.
@@ 10 Complexity Analysis of CARA Impact
During decoding, LLaVA’s per‑step attention scales with the whole prefix $(T+L)$, yielding a total cost $\mathcal{O}(nA(T+L)+nA^{2})$ for $A$ generated tokens. In our model, self‑attention is unchanged while cross‑attention is linear in $M$ and independent of $A$ on the key/value side, giving $\boxed{\mathcal{O}(n(AT+A^{2})+n_{\times}AM)}$ . Memory usage follows the same trend: LLaVA’s KV cache grows as $\mathcal{O}(nd(T+L+A))$, whereas ours is $\mathcal{O}(nd(T+A))+\mathcal{O}(n_{\times}Md)$ with a constant visual term. The gains are most pronounced when $L$ is large (e.g., WSI), since we replace the $\mathcal{O}(L)$ dependence with $\mathcal{O}(M)$ both in prefill compute and in decoding’s per‑step cost.

### v2

@@ 3 Methodology
A WSI is decomposed into a long sequence of tiles. For each training image-text pair, we denote the tile-level tokens from the tile-level encoder by $\bm{X}=[\bm{x}_{1},\ldots,\bm{x}_{N}]\in\mathbb{R}^{N\times D_{v}}$, with 2D-grid coordinates $\bm{C}=(c_{i},r_{i})\in\mathbb{N}^{2}$, where NN is the number of the tile-level tokens, $D_{v}$ is the tile-level encoder’s feature dimension. $c_{i}$ and $r_{i}$ are the column and row grid coordinates with stride 11, respectively. In MAE pre-training task, we produce $L$-length visual latents $\bm{Z}\in\mathbb{R}^{L\times D_{v}^{\prime}}$, where $D_{v}^{\prime}$ is the embedding dimension of MAE resampler. The text query tokens are $\bm{Q}_{\text{text}}$ and the answer labels are $\bm{Y}$. In single-round visual question answering (VQA) tasks, $\bm{\hat{Y}}$ is generated by an MLLM for a given $\bm{X}$ and text query $\bm{Q}_{\text{text}}$.
@@ 6 Redundancy and Task Relevance Analysis
$F(p)=\frac{\sum_{i=1}^{\lfloor pN\rfloor}\max(s_{i},0)}{\sum_{i=1}^{N}\max(s_{i},0)},\quad p\in[0,1].$
@@ 7 Pairwise Ranking Loss Details
To compute this loss, we first sort the Top-$M$ indices II according to the teacher distribution $\bm{q}$, yielding $I_{\text{sorted}}$. We select $n=\min(M/2,N_{\text{pairs}})$ pairs, where $N_{\text{pairs}}$ is a hyperparameter (e.g., 8). We define the ”high” set $I_{\mathrm{hi}}=I_{\text{sorted}}[:n]$ (the top $n$ indices) and the ”low” set $I_{\mathrm{lo}}=I_{\text{sorted}}[M//2:M//2+n]$ (the first $n$ indices from the bottom half).
@@ 8 Implementation and Training Details
Training Details. The resampler is pre-trained for $100\mathrm{k}$ iterations on CONCH-V1 [32] WSI features with a mini-batch size of 22 per GPU and gradient accumulation 44 on 44 RTX A6000 Ada GPUs (effective global batch size $32$). We randomly mask $75\%$ of the input tokens and optimize the MAE objective $\mathcal{L}_{\mathrm{MAE}}$ introduced in Sec. 3.1, with $\lambda_{\mathrm{cover}}=5\times 10^{-4}$ for the coverage term $\mathcal{L}_{\mathrm{cover}}$ and $\lambda_{\mathrm{feat}}=1\times 10^{-3}$ for the diversity term $\mathcal{L}_{\mathrm{feat}}$. We use AdamW with an initial learning rate $1\times 10^{-4}$, a $3\%$ warm-up phase, and a cosine learning-rate schedule. During this pre-training we gradually anneal the Top-$K$ in the resampler cross-attention from $(128,64)$ to $(64,32)$ following a cosine schedule.
We then train the first fusion stage for 33 epochs with a global batch size of $512$ (batch size 22 per GPU and gradient accumulation $64$ on 44 GPUs) and an initial learning rate $1\times 10^{-4}$. In this stage we freeze both the Qwen2.5-7B-Instruct LLM and the resampler, and only train the visual projector together with the CARA+TIS modules on top of the LLM. The training objective is the multimodal loss as defined in Sec. 3.2, where we set $\lambda_{\mathrm{TIS}}=\lambda_{\mathrm{rank}}=0.02$ for the self-distillation (TIS) loss $\mathcal{L}_{\mathrm{TIS}}$ and the ranking loss $\mathcal{L}_{\mathrm{rank}}$.
In the second fusion stage, we unfreeze the resampler’s last cross-attention layer and its latent in-projection, and insert LoRA modules (rank $128$, $\alpha=64$) into all linear layers of the LLM. We then train all unfrozen components on all VQA pairs in WSI-Bench for 22 epochs with a global batch size of $128$ and an initial learning rate $2\times 10^{-5}$, again using the same multimodal fusion loss with $\lambda_{\mathrm{TIS}}=\lambda_{\mathrm{rank}}=0.02$. Across all training stages, we use Xtuner [10] with DeepSpeed ZeRO-2 [40] and FlashAttention 2 [11] for efficient training; FA2 is applied to the LLM self-attention but not to the cross-attention layers. A cosine learning-rate scheduler is used in all stages, and all experiments are conducted on 44 RTX A6000 Ada GPUs.
Training Optimization. The pre-training requires about $10$ hours. The first fusion stage takes about 11 hour. The second fusion stage requires about 88 hours to 99 hours.
@@
With file format conversion, we are able to load long sequences with very low I/O costs. Unlike the conventional tile-level data loader in digital pathology [9, 51], loading one data point in a single batch, we are able to load multiple tile-level sequences in a single batch simultaneously. To deal with sequences with different lengths, we pack these sequences into the same length (we pad sequences with different lengths to the same $60k$ lengths and load padded sequences with their corresponding padding masks). As a result, we can increase the batch size to more than one (e.g 4) on each GPU. Doing so allows us to reduce the data loading rounds if we set the max training epochs to be 22.
@@ 10 Complexity Analysis of CARA Impact
During decoding, LLaVA’s per‑step attention scales with the whole prefix $(T+L)$, yielding a total cost $\mathcal{O}(nA(T+L)+nA^{2})$ for AA generated tokens. In our model, self‑attention is unchanged while cross‑attention is linear in $M$ and independent of AA on the key/value side, giving $\boxed{\mathcal{O}(n(AT+A^{2})+n_{\times}AM)}$ . Memory usage follows the same trend: LLaVA’s KV cache grows as $\mathcal{O}(nd(T+L+A))$, whereas ours is $\mathcal{O}(nd(T+A))+\mathcal{O}(n_{\times}Md)$ with a constant visual term. The gains are most pronounced when $L$ is large (e.g., WSI), since we replace the $\mathcal{O}(L)$ dependence with $\mathcal{O}(M)$ both in prefill compute and in decoding’s per‑step cost.